# Electroweak Monopole-Antimonopole Pair Production at LHC

### v2

@@ III Monopole Production in Early Universe: A Review
To amplify this point we start from the temperature dependent effective action of the standard model expressed by [27, 38, 39, 40, 41, 42, 43]
@@
where $V_{0}$ (with $\lambda\simeq 0.26$ and $\rho_{0}\simeq 246~{}{\rm GeV}$) is the zero-temperature potential, $N_{*}$ is the total number of distinct helicity states of the particles with mass smaller than $T$ (counting fermions with the factor 7/8), $C_{1}$ and $C_{2}$ are the contributions from the $\rho$ dependent part of the masses of gauge bosons, Higgs field, and fermions, $m_{t}$ is the top quark mass, and $\delta V_{T}$ is the slow-varying logarithmic corrections and the lighter quark contributions which we will neglect from now on.
@@
At $T_{c}$ the new vacuum bubbles start to nucleate at $\rho_{s}=0$, which tunnels to the stable vacuum $\rho_{+}$ after $T_{c}$. Below this critical temperature $\rho_{+}$ becomes the true minimum of the effective potential, but $\rho_{s}=0$ remains a local (unstable) minimum till the temperature reaches $T_{2}$,
@@
Figure 3: The effective potential (21) at different temperatures. Notice that the potentials at $T_{1},~{}T_{c},~{}T_{2}$ are almost indistinguishable. Here the unit of $V_{eff}$ is chosen to be $V_{0}=(\lambda/8)\rho_{0}^{4}=1$.
The monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism, so that the monopole production start from $T_{c}$. However, we have to keep in mind that the thermal fluctuations of the Higgs vacuum which create the seed of the monopoles continue as long as [27, 37]
@@
To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [44]
@@
=$=\left\{\begin{array}[]{ll}\dfrac{1}{2}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}M_{H}^{2},&~{}~{}~{}T>T_{c},\\ \dfrac{1}{2}\Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}M_{H}^{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
@@
Figure 5: The temperature dependent Higgs and W-boson masses. The blue and red curves represent the Higgs and W-boson masses.
$\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature. But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$. Moreover, we have
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$). This tells that the baby monopole mass can be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value. This could allow LHC to produce the monopole even when the adolescent mass becomes bigger than 7 TeV. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
At this point one might wonder if we could also observe the Higgs and W bosons which have the temperature dependent masses (the “baby” Higgs and W bosons) at LHC. This is not possible, because they are ordinary particles produced by the ordinary process, not the topological particles produced topologically. So there is no reason to believe that LHC can produce them.
@@
IV Electroweak monopoles from a fireball at LHC
If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it is commonly believed, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does. In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
@@ IV Electroweak monopoles from a fireball at LHC
The above discussion tells that LHC could produce the electroweak baby monopole which could have the mass considerably less than the adolescent mass of 11.0 TeV, if we assume that the monopole production mechanism at LHC is topological. This is unexpected.
If so, can we really detect the infant electroweak monopole at LHC whose mass is smaller than the adolescent mass? To answer this we have to keep in mind that in the early universe most of the monopole-antimonopole pairs are quickly annhilated as soon as they are produced. This is because at the production temperature the capture radius of the monopole and anti-monopole is almost hundred times bigger than the correlation length. As the result most of the monopole-antimonopole pairs are annihilated as soon as they are produced, and only about $10^{-5}$ of the monopoles are survived in the early universe [27]. So one might wonder if we could have the same annihilation of the monopole-antimonopole pairs at LHC.
Actually, this annhilation is not expected to play an important role at LHC. This is because (unlike the early universe) the size of the p-p fireball at LHC is only about thirteen times bigger than the correlation volume, so that many of the monopoles could quickly escape the fireball before they are annhilated. This must be clear from (45) and (46). Moreover, once they escape from the fireball they are quickly freezed out (i.e., “quenched”). And as far as we know, nothing can destroy these quenched monopoles after they escape from the fireball. In this case MoEDAL should be able to detect these infant monopoles in principle.
On the other hand, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball. This necessiate us to study the monopolium bound state which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
Let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
@@
where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 11.0~{}\text{TeV}$. So the Rydburg energy of the monopolium becomes 51.615 TeV. Moreover, the Bohr radius $R$ of the monopolium is given by
@@
$\simeq 1.4\times 10^{-24}~{}\mbox{cm},$ (51)
Notice, however, that these results are totally unreasonable. First, the above Rydburg energy is too big (roughly 2,346 times bigger than two monopole masses), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium. Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
This could mean the followings. First, the monopolium quickly annihilates itself, long before it reaches to the ground state. In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy. Or else, this could mean that the naive atomic description of the monopolium is not acceptable. In this case we have to find a more realistic model of the monopolium.
The reason for the above results originates from two facts. First, the magnetic coupling of the monopole is too strong. Second, the monopole mass is too big. We could make the atomic model more realistic by improving on these points. For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$. Moreover, we could make the monopole mass less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
On the other hand, the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV. This could make the monopolium mass considerably lighter than $2M$, so that we could have the monopolium of mass less than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV. With these modifications, we could have a more realistic model of the electroweak monopolium that we discuss now.
VI Electroweak Monopolium: A Realistic Model
To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, because the monopole (and antimonopole) has the size of $1/M_{W}$. This means that at short distances the singularity of the magnetic potential in (49) should be regularized.
@@ VI Electroweak Monopolium: A Realistic Model
$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ \\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
@@
Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium is shown as a function of $r/r_{0}$.
@@
$a_{\rm I}=\frac{2l+3+x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4},~{}~{}~{}~{}b_{\rm I}=l+3/2,$
@@
as $x$ goes to infinity. Hence, the general solution in the second region has a following form,
@@
$\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$ (74)
$+2\sqrt{x_{0}}\big{(}(1+l)\sqrt{\epsilon}-1\big{)}\frac{U\Big{(}\begin{matrix}2+l-\epsilon^{-1/2}\\ 3+2l\end{matrix}\Big{|}2x_{0}\epsilon^{1/2}\Big{)}}{U\Big{(}\begin{matrix}1+l-\epsilon^{-1/2}\\ 2+2l\end{matrix}\Bigr{|}2x_{0}\epsilon^{1/2}\Big{)}}$ (79)
@@
Figure 8: The contours where Eq. (8) holds as function of $\sqrt{x_{0}}$ and $\sqrt{\epsilon}$ with $l=0$. The dashed lines represent harmonic oscillator levels given by (82). The horizontal dotted lines are the levels of the hydrogen atom, i.e., $\epsilon=1/n^{2}$, with $n=1,2,\ldots$
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (80). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 for $l=0$.
@@
Notice that in the above analysis we have assumed that the monopole mass is 11 TeV. However, at LHC the monopole mass at $T_{min}^{LHC}$ becomes 5.3 TeV, so that the monopolium at LHC is expected to be the monopole-antimonopole bound state with mass 10.6 TeV. This will change the energy spectrum and the Bohr radius of the monopolium. For example, $E_{0}$ in (84) should change to $E_{0}\simeq-8.1~{}\text{TeV}$, which reduces the monopolium mass to 2.5 TeV at LHC.
It should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential
@@
has been discussed before [33]. Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar. In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$. This implies that both could be used as a realistic model of the monopolium.

### v4

@@ III Monopole Production in Early Universe: A Review
To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak monopole production in the early universe [15, 25, 26, 18]
@@
where $V_{0}$ (with $\lambda\simeq 0.26$ and $\rho_{0}\simeq 246~{}{\rm GeV}$) is the zero-temperature potential, $C_{1}$ and $C_{2}$ terms are the loop contributions from the gauge bosons, Higgs field, and heavy fermions, $N_{*}$ is the total number of distinct helicity states of the particles with mass smaller than $T$ (counting fermions with the factor 7/8), $m_{t}$ is the top quark mass, and $\delta V_{T}$ is the slow-varying logarithmic corrections and the lighter quark contributions which we will neglect from now on.
@@
Figure 3: The effective potential (21) at different temperatures. Notice that the potential at $T_{1},~{}T_{c},~{}T_{2}$ are almost indistinguishable. Here the unit of $V_{eff}$ is chosen to be $V_{0}=(\lambda/8)\rho_{0}^{4}=1$.
@@
At $T_{c}$ the new vacuum bubbles start to nucleate at $\rho_{s}=0$, which tunnels to the stable vacuum $\rho_{+}$ after $T_{c}$. Below this critical temperature $\rho_{+}$ becomes the true minimum of the effective potential, but $\rho_{s}=0$ remains a local minimum till the temperature reaches $T_{2}$,
@@
The monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism, so that the monopole production start from $T_{c}$. And the thermal fluctuations of the Higgs vacuum which create the seed of the monopoles continue as long as we have [15, 24]
@@
To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [kolb]
@@
Figure 5: The temperature dependent Higgs and W boson masses. The blue and red curves represent the Higgs and W boson masses.
@@
=$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
@@
$\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature. But the W-boson which is massless above $T_{c}$ (before the symmetry breaking) becomes massive at $T_{c}$, and we have
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$). This confirms that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value. This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
IV Electroweak Monopole Production from p-p Fireball at LHC
If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does. In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
@@ IV Electroweak Monopole Production from p-p Fireball at LHC
The above discussion tells that LHC could produce the electroweak baby monopole of mass around 5.3 TeV even when the adolescent mass becomes 11.0 TeV, if we assume that the monopole production mechanism at LHC is thermal. If this is true, this would be really remarkable.
Unfortunately, the above argument has a critical defect, because here we have implicitly assumed that the temperature of the p-p fireball would be high enough to produce the monopole thermally (assuming that the 14 TeV p-p fireball energy would produce the high temperature necessary for LHC to produce the monopole thermally). But this is not obvious, because the temperature and energy of the fireball are two different things. As we have argued, for LHC to produce the monopole thermally, the temperature of the fireball must be no less than 119.4 GeV. And if the temperature of the fireball higher than this, the above conclusion becomes valid. But is this really so? How can we guarantee this? This is a non-trivial question.
A straightforward way to answer this question is to measure the temperature of the p-p firebal at LHC. And actually the ALICE group has measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [27]. The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature to produce the monopole thermally, even with the future FCC energy upgrade.
This tells that here again, the thermal production of the electroweak monopole at LHC is practically impossible, although it is logically possible. This teaches us an important lesson. If the monopole production mechanism at LHC is the thermal production, it is vertually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low. In this case the only way to detect the electroweak monopole is to look for the remnant monopoles produced in the earluy universe.
This also tells that the only probable way for LHC to prooduce the monopole is by the Drell-Yan process and/or the Schwinger mechanism. Obviously, the present 14 TeV LHC could produce the monopole if the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism. But as we have pointed out, in this case there is a possibility for LHC to produce the monopole in the form of the monopolium bound states, even when the monopole mass becomes 11 TeV. This is because the monopolium binding energy could reduse the monopolium mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV. This necessiate us to study the monopolium bound state, which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
The monopolium itself has been studied before for various reasons [28, 29, 30]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV [28]. But here we are interested in the electroweak monopolium made of the Cho-Maison monopole pair.
To discuss the electroweak monopolium we let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
@@
where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 11.0~{}\text{TeV}$. So the Rydburg energy of the monopolium becomes 51,615 TeV. Moreover, the Bohr radius $R$ of the monopolium is given by
@@
$\simeq 1.4\times 10^{-24}cm,$ (51)
Notice, however, that these results are totally unreasonable. First, the above Rydburg energy is too big (roughly 2,346 times bigger than the monopole-antimonopole mass), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium. Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
This could mean the following. First, the monopolium quickly annihilates itself, long before it reaches to the ground state. In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy. Or else, this could mean that the naive atomic description of the monopolium is not acceptable. In this case we have to find a more realistic model of the monopolium.
The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monipolium size too small. We could make the atomic model of the monopolium more realistic by improving on this point. For example, we could modify the magnetic Coulombic potential at short distance for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
On the other hand the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV. This could make the monopolium mass considerably less than $2M$, so that we could have the monopolium of mass lighter than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV, with a more realistic monopolium potential.
VI A Realistic Electroweak Monopolium
To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole (and antiminopole) has the size of $1/M_{W}$. This means that at short distance the singularity of the magnetic potential in (49) should be regularized.
@@ VI A Realistic Electroweak Monopolium
$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ {}\hfil\\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},{}{}{}{}{}{}{}{}&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
@@
Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium shown as a function of $r/r_{0}$.
@@
$a_{\rm I}=\frac{2l+3+x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4},~{}~{}~{}~{}b_{\rm I}=l+3/2,$ (60)
and
@@
as $x$ goes to infinity. Hence, the general solution in the second region has the following form,
@@
Figure 8: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$. The dashed lines represent harmonic oscillator energy levels given by (67). The horizontal dotted lines are the energy levels of the hydrogen atom given by (66).
@@
$\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$
$+2\sqrt{x_{0}}\big{(}(1+l)\sqrt{\epsilon}-1\big{)}\frac{U\Big{(}\begin{matrix}2+l-\epsilon^{-1/2}\\ 3+2l\end{matrix}\Big{|}2x_{0}\epsilon^{1/2}\Big{)}}{U\Big{(}\begin{matrix}1+l-\epsilon^{-1/2}\\ 2+2l\end{matrix}\Bigr{|}2x_{0}\epsilon^{1/2}\Big{)}}$
@@
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 (for $l=0$).
@@
Notice that in the above analysis we have assumed the monopole mass to be 11.0 TeV to have the monopolium mass 5.2 TeV. And (69) tells that the monopolium mass becomes 3.3 TeV when the monopole mass becomes 7 TeV. Perhaps more importantly, this tells that the maximum mass of the monopole that the present 14 TeV LHC can produce the monopolium of mass 14 TeV becomes 29.8 TeV. This strongly implies that the present LHC could produce the electroweak monopole in the form of the monopolium bound states for any reasonable monopole mass (as far as the mass remains less than 29.8 TeV), if the monopole production mechanisn becomes the Drell-Yan process.
Of course, the monopolium mass at LHC may not turn out to be the same as the above prediction, because we do not know how realistic our potential (55) is. In this connection it should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential
@@
has been discussed before [30]. Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar. In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$. This implies that our results discussed above are reliable.